# Zaragoza, Spain Feature Splatting

### v1

@@ A.1 Implementation
We use a two layer MLP with bias. The first layer has 64 neurons and a SiLU activation function [Hendrycks and Gimpel(2023)]. For the only RGB implementation, the output layers have 3 output neurons and a Sigmoid activation function. Instead, for the semantic segmentation implementation, the output layer has 67 output neurons, with a sigmoid activation function for the first 3 channels, and a Softmax for the last 64. We use the Adam optimizer with $\epsilon=1\times 10^{-15}$, a learning rate of $0.001$ for the MLP and $0.0025$ for the feature vectors, and $\lambda_{sem}=0.001$ for the semantic segmentation model. The rest of the Gaussians’ parameters and optimization hyperparameters follow 3DGS’s [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] original implementation. We use the 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] CUDA differentiable rasterizer. We adapt it to our model by skipping spherical harmonics and passing feature vectors as precomputed colors. To do it, we change the number of channels when compiling from 3 to 16 and 32 for FeatSplat-16 and FeatSplat-32 respectively. The model has been implemented on Pytorch 1.12 and CUDA SDK 11.6.2.
@@ A.2 Embeddings ablation
The ablation results in Table 3 show that both $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ have by themselves a positive impact on the metrics, while $\mathbf{x}_{cam}$ does not improve against not adding it. While $\mathbf{e}_{p}$ is responsible from almost all the improvement on Train, $\mathbf{x}_{cam}$ has a better performance on Bicycle, and both together achieve the best performance. On the other hand, $\mathbf{x}_{rot}$ alone does not improve the performance. Combining it with $\mathbf{x}_{cam}$ improves the performance of $\mathbf{x}_{cam}$ alone on Train, although less than $\mathbf{e}_{p}$. The results show that $\mathbf{x}_{cam}$ is redundant, with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ having the best performance.
Appendix B Datasets
B.1 Mip360, T&T and DB.
The three typical datasets we use amount to a total of 12 scenes: Mip360[Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], with 9 scenes, 5 outdoor at $1237\times 822$ and 4 indoor at $1558\times 1039$, with an average of 216 views per scene; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with 2 outdoor scenes at $979\times 546$ and an average of 276 views per scene; and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] with 2 indoor scenes at $1264\times 832$ and an average of 244 views per scene. In these datasets the testing set is selected sampling one every eight frames of the camera trajectory, using the other 7 frames to train.
B.2 ScanNet++
@@
B.3 Test and Train cameras plots
Figures 879 and 10 show the distribution of train and test cameras for all evaluated scenes.

### v2

@@ A.1 Implementation
We use a two-layer MLP with bias. The first layer has 64 neurons and a SiLU activation function [Hendrycks and Gimpel(2023)]. For the RGB-only implementation, the output layers have 3 output neurons and a Sigmoid activation function. Instead, for the semantic segmentation implementation, the output layer has 67 output neurons, with a Sigmoid activation function for the first 3 channels, and a Softmax for the last 64. We use the Adam optimizer with $\epsilon=1\times 10^{-15}$, a learning rate of $0.001$ for the MLP and $0.0025$ for the feature vectors, and $\lambda_{sem}=0.001$ for the semantic segmentation model. The rest of the Gaussians’ parameters and optimization hyperparameters follow 3DGS’s [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] original implementation. We use the 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] CUDA differentiable rasterizer. We adapt it to our model by skipping spherical harmonics and passing feature vectors as precomputed colors. To do it, we change the number of channels when compiling from 3 to 16 and 32 for FeatSplat-16 and FeatSplat-32 respectively. The model has been implemented on Pytorch 1.12 and CUDA SDK 11.6.2.
@@ A.2 Embeddings ablation
The ablation results in Table 3 show that both $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ have by themselves a positive impact on the metrics, while $\mathbf{x}_{cam}$ does not improve against not adding it. While $\mathbf{e}_{p}$ is responsible from almost all the improvement on Train, $\mathbf{x}_{cam}$ has a better performance on Bicycle, and both together achieve the best performance. On the other hand, $\mathbf{x}_{rot}$ alone does not improve the performance. Combining it with $\mathbf{x}_{cam}$ improves the performance of $\mathbf{x}_{cam}$ alone on Train, although less than $\mathbf{e}_{p}$. The results show that $\mathbf{x}_{rot}$ is redundant, with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ having the best performance.
Appendix B ScanNet++ Novel View Synthesis Benchmark
ScanNet++
SSIM PSNR LPIPS FPS
HW GTS - RPBG 0.873 24.355 0.280 $\thicksim 1$
FeatSplat–32 (Ours) 0.869 24.247 0.314 50
Nerfacto 0.861 24.049 0.342 0.3
TensoRF 0.849 23.978 0.407
3D Gaussian Splatting 0.871 23.891 0.319 155
iNGP 0.859 23.812 0.375 0.3
Table 4: ScanNet++ Novel View Synthesis Benchmark
Appendix C Datasets
C.1 Mip360, T&T and DB.
The three typical datasets we use amount to a total of 12 scenes: Mip360[Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], with 9 scenes, 5 outdoor at $1237\times 822$ (Bicycle, Flowers, Garden, Stump, Treehill,) and 4 indoor at $1558\times 1039$ (Bonsai, Counter, Kitchen, Room), with an average of 216 views per scene; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with 2 outdoor scenes at $979\times 546$ (Train and Truck) and an average of 276 views per scene; and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] with 2 indoor scenes (DrJohnson and Playroom) at $1264\times 832$ and an average of 244 views per scene. In these datasets the testing set is selected sampling one every eight frames of the camera trajectory, using the other 7 frames to train.
C.2 ScanNet++
@@
C.3 Test and Train cameras plots
Figures 9810 and 11 show the distribution of train and test cameras for all evaluated scenes.